# pMixFed: Efficient Personalized Federated Learning through Adaptive Layer-Wise Mixup

### v1

@@ 3 Problem Formulation
where $|\mathcal{D}|=\sum_{k=1}^{K}|\mathcal{D}_{k}|$ and $F(\theta)$ is the global loss function of FedAvg [32]. FL is performed in at iterative fashion. At each round, each client downloads the current version of the global model and trains it using their local data. Clients then send the updated model parameters to the central server. The central server uses the model updates from the selected clients and aggregates them to update the global model. Iterations continue until convergence.
In Equation 1, the assumption is that the data is collected from an IID distribution, and all clients should train their data according to the exact same model. However, this assumption cannot be applied within many practical FL settings due to the non-IID nature of data and resource limitations [18, 17]. In the PFL settings with high heterogeneity and non-IID data distribution, the same issue persists and the local parameters need to be customized toward each agent. PFL extends FL by solving the following [26, 34]:
@@
PFL explicitly handles data heterogeneity through the term $\mathcal{F}_{k}(\theta_{k})$ which accounts for model heterogeneity by considering personalized parameter $\theta_{k}$ for client $k$. Meanwhile, $\theta$ represents the shared global model parameters in Equation 2, and $\alpha_{k}$ acts as a regularizer and indicates the degree of personalization tuning collaborative learning between personalized local models $\theta_{k}$ and generalized global model $\theta$. When $\alpha_{k}$ is small, the personalization power of the local models will be increased. If $\alpha_{k}$ is large, the local models’ parameters tend to be closer to the global model parameters $\theta$.
@@ 3.1 Partial Personalized Model
Among different partitioning strategies for partial PFL [34], the most popular technique is to assign local personalized layers $L_{l,k}^{t}$ to final layers and allow the base layers $L_{g,k}^{t}$ to share the knowledge similar to FedPer [4]. This choice aligns with insights from MAML  [^3] Model-Agnostic Meta-Learning algorithm, suggesting that initial layers keeps general and broad information while personalized characteristics manifest prominently in the higher layers. Accordingly, we would have:
@@
For simplicity $L_{g}=L_{g,k}$ and $L_{l}=L_{g,k}$ where $\{1\leq g\leq s\leq l\leq M\}$ and $s$ is the split(cut) layer. The objective in solving Equation 3.1 is to find the optimal $s$ (cut layer) which minimizes the personalization objective: $\sum_{k=1}^{K}\frac{1}{|\mathcal{D}_{k}|}\mathcal{F}_{k}(\theta_{k})=L_{l}^{(t)}(L_{g}^{(t)}(x_{k}))$.
@@
In partial models, after several rounds of local training, both personalized and global layers of local model are updated. This update could be synchronous like FedSim or asynchronous as in FedAlt. The Personalized layers will be frozen until the next communication round, $L_{l}^{(t+1)}=L^{\prime}_{l}$ and the global layers will be sent to the server for global model aggregation : $G^{(t+1)}\leftarrow\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}L^{\prime}_{g}(x_{k})$. In the next broadcasting phase, the shared layers of the local model will be updated as $L_{g}^{(t+1)}\leftarrow G^{(t+1)}$. Our goal is to benefit from Mixup to improve personalization.
4 The Proposed Method
@@ 4.1 Mixup
Mixup is a data augmentation technique for enhancing model generalization [61] based on learning to generalize on linear combinations of training examples. Variations of Mixup have consistently excelled in vision tasks, contributing to improved robustness, generalization, and adversarial privacy. Mixup creates augmented samples as:
$\small\begin{split}\bar{x}=\lambda.x_{i}+(1-\lambda).x_{j}\\ \bar{y}=\lambda.y_{i}+(1-\lambda).y_{j},\end{split}$ (4)
where $x_{i}$ and $x_{j}$ are two input samples, $y_{i}$ and $y_{j}$ are the corresponding labels, $\lambda$, $\lambda\sim\text{Beta}(\alpha,\alpha),\lambda\in[0,1]$, is the degree of interpolation between the two samples. Mixup relates data points belonging to different classes which has been shown to be successful in mitigating overfitting and improving model generalization [49, 13, 62]. There are many variants of Mixup that has been developed to address specific challenges and enhance its effectiveness. For instance, AlignMixup [48] improves local spatial alignment by introducing transformations that better preserve semantic consistency between input pairs. Manifold Mixup [49] extends the concept to hidden layer representations, acting as a powerful regularization technique by training deep neural networks (DNNs) on linear combinations of intermediate features. CutMix [60] replaces patches between images, blending visual information while retaining spatial structure. Remix [7] addresses class imbalance by assigning higher weights to minority classes during the mixing process, enhancing the robustness of the trained model on imbalanced datasets. AdaMix [13] dynamically optimizes the mixing distributions, reducing overlaps and improving training efficiency. This linear interpolation also serves as a regularization technique that shapes smoother decision boundaries, thereby enhancing the ability of a trained model to generalize to unseen data. Mixup can also increase robustness against adversarial attacks [62, 5]; and improves performance against noise, corrupted labels, and uncertainty as it relaxes the dependency on specific information [13].
Figure 2: Workflow of pMixFed: Mixup is used in two stages. 1-Broadcasting: when transferring knowledge to local models, the frozen personalized model $L_{k}^{(t)}$ is mixed up with global model $G^{(t)}$ according to the adaptive mix factor $\mu_{k}^{(t)}$ which determines layer-wise mixup degree $\lambda_{i}$ for layer $i$. 2-Aggregation: The updated global model $G^{(t+1)}$ is generated through a Mixup between the updated local model $L^{(t+1)}$ and the current shared global model $G^{\prime(t)}$. Consequently, a state of the previous global model $G^{\prime(t)}$ is retained, which aids in mitigating catastrophic forgetting.
@@
Our goal is to leverage the well-established benefits of Mixup in the context of personalized FL. While Mixup has previously been employed in FL frameworks, such as XORMixup [40], FEDMIX [57], and FedMix [51], prior studies have primarily focused on using Mixup for data augmentation or data averaging. We propose pMixFed by integrating Mixup on the model parameter space, rather using it on the feature space. We apply Mixup between the parameters of the global and the local models in a layer-wise manner for more customized and adaptive PFL. Our approach eliminates the need for static and rigid partitioning strategies. Specifically, during both the broadcasting and aggregation stages of our partial PFL framework, we generate mixed model weights using a an interpolation strategy which is illustrated in Figure 2. Mixup offers the flexibility in combining models by introducing a mix degree for each layer $\lambda_{i}$, which changes gradually according to $\mu$, i.e., the mix factor. Parameter $\mu$ is also updated adaptively in each communication round and for each client according to the test accuracy of the global and the local model during the evaluation phase of FL. $\mu$ is computed as follows:
$\small\mu_{k}^{t}=1-\frac{1}{1+e^{-\delta(Acc^{b}-1)}}$ (5)
$\delta=(t/T)$, where $t$, and $T$ are the current communication round, and the overall number of communication rounds respectively. Also, $Acc$ calculated as: $Acc=(Acc_{k}^{t}/Acc_{overall}^{(t-1)})$ in the broadcasting phase and average test accuracy of the previous global model $\left(Acc=G^{t}(x,\theta^{t}),y\right)$ on all local test sets $x=\{x_{1},x_{2},\ldots,x_{K}\}$. , in the aggregation stage. $b$ is the offset parameter for sigmoid function which is set to $2$ after several experiments. More detailed discussion on parameter $\mu$’s rule of update is discussed in section 5 and the reason behind this formulation is discussed in the Appendix 8.
@@
$\small\begin{split}\text{Broadcasting Stage:}\quad\quad\lambda_{i}=\begin{cases}1&\lambda_{i}>1\\ \mu*(n-i)&\lambda_{i}\leq 1\end{cases}\\ \text{Aggregation Stage:}\quad\quad\lambda_{i}=\begin{cases}0&\lambda_{i}\leq 0\\ 1-(i*\mu)&\lambda_{i}>0,\end{cases}\end{split}$ (6)

### v2

@@ 3 Problem Formulation
where $|\mathcal{D}|=\sum_{k=1}^{K}|\mathcal{D}_{k}|$ and $F(\theta)$ is the global loss function of FedAvg mcmahan2017communication. FL is performed in at iterative fashion. At each round, each client downloads the current version of the global model and trains it using their local data. Clients then send the updated model parameters to the central server. The central server uses the model updates from the selected clients and aggregates them to update the global model. Iterations continue until convergence.
In Equation 1, the assumption is that the data is collected from an IID distribution, and all clients should train their data according to the exact same model. However, this assumption cannot be applied within many practical FL settings due to the non-IID nature of data and resource limitations imteaj2021survey; imteaj2021fedparl. In the PFL settings with high heterogeneity and non-IID data distribution, the same issue persists and the local parameters need to be customized toward each agent. PFL extends FL by solving the following li2021model; pillutla2022federated:
@@
PFL explicitly handles data heterogeneity through the term $\mathcal{F}_{k}(\theta_{k})$ which accounts for model heterogeneity by considering personalized parameter $\theta_{k}$ for client $k$. Meanwhile, $\theta$ represents the shared global model parameters in Equation 2, and $\alpha_{k}$ acts as a regularizer and indicates the degree of personalization tuning collaborative learning between personalized local models $\theta_{k}$ and generalized global model $\theta$. When $\alpha_{k}$ is small, the personalization power of the local models will be increased. If $\alpha_{k}$ is large, the local models’ parameters tend to be closer to the global parameters $\theta$.
@@ 3.1 Partial Personalized Model
Among different partitioning strategies for partial PFL pillutla2022federated, the most popular technique is to assign local personalized layers $L_{l,k}^{t}$ to final layers and allow the base layers $L_{g,k}^{t}$ to share the knowledge similar to FedPer arivazhagan2019federated. This choice aligns with insights from MAML  [^3] Model-Agnostic Meta-Learning algorithm, suggesting that initial layers keeps general and broad information while personalized characteristics manifest prominently in the higher layers. Accordingly, we would have:
@@
For simplicity $L_{g}=L_{g,k}$ and $L_{l}=L_{g,k}$ where $\{1\leq g\leq s\leq l\leq M\}$ and $s$ is the split(cut) layer. The objective in solving Equation 3.1 is to find the optimal $s$ (cut layer) which minimizes the personalization objective: $\sum_{k=1}^{K}\frac{1}{|\mathcal{D}_{k}|}\mathcal{F}_{k}(\theta_{k})=L_{l}^{(t)}(L_{g}^{(t)}(x_{k}))$. In partial models, after several rounds of local training, both personalized and global layers of local model are updated. This update could be synchronous like FedSim or asynchronous as in FedAlt. The Personalized layers will be frozen until the next communication round, $L_{l}^{(t+1)}=L^{\prime}_{l}$ and the global layers will be sent to the server for global model aggregation : $G^{(t+1)}\leftarrow\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}L^{\prime}_{g}(x_{k})$. In the next broadcasting phase, the shared layers of the local model will be updated as $L_{g}^{(t+1)}\leftarrow G^{(t+1)}$.
Table 1: Notation used throughout the paper. Vectors are column vectors; $\|\cdot\|$ is the Euclidean norm.
Symbol Type Meaning
$K$ scalar Number of clients
$\mathcal{U}_{t}$ set Clients selected at round $t$; $|\mathcal{U}_{t}|{=}K$
$D_{k}$ dataset Local dataset of client $k$; size $|D_{k}|$
$|D|$ scalar Total data size, $|D|{=}\sum_{k}|D_{k}|$
$\Omega_{k}$ scalar Aggregation weight $|D_{k}|/|D|$
$\ell(\cdot,\cdot)$ function Pointwise loss used to build local/global objectives
$f_{k}(\theta,x)$ function Model prediction on client $k$ with parameters $\theta$
$F_{k}(\theta)$ function Local objective on client $k$
$F(\theta)$ function Global objective $\sum_{k}\Omega_{k}F_{k}(\theta)$
$\theta$ vector Global (server) model parameters
$\theta_{k}$ vector Client $k$’s local/personal parameters (if stored)
$G^{(t)}$ vector Server/global parameters at round $t$ (alias of $\theta^{t}$)
$L_{k}^{(t)}$ vector Local model of client $k$ at round $t$
$L^{(t)}_{g,k}$ vector Global/shared layers of $L_{k}^{(t)}$
$L^{(t)}_{l,k}$ vector Local/personalized layers of $L_{k}^{(t)}$
$M$ scalar Number of layers in the network
$s$ index Split (cut) layer index, $1\leq s\leq M$
$t$ index Communication round index ($t{=}0,\dots,T{-}1$)
$T$ scalar Total number of communication rounds
$r,\ \tau$ scalar Number of local steps per round (local iterations/epochs)
$b$ scalar Local mini-batch size
$\eta_{\ell}$ scalar Local learning rate (client step size)
$\eta_{g}$ scalar Server/aggregation step size in the FedSGD view
$\lambda^{t}_{k,i}$ scalar Mixup coefficient for client $k$, layer $i$, round $t$ ($\in[0,1]$)
$\lambda^{t}_{k}$ scalar Client-level mixup (uniform across layers if assumed)
$\bar{\lambda}^{\,t}$ scalar Population-averaged mixup, $\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\lambda_{k}^{t}$
$\eta_{\mathrm{eff}}^{\,t}$ scalar Effective step size $(1{-}\bar{\lambda}^{\,t})\,\eta_{\ell}$ in SGD view
$\mu$ scalar Mix factor controlling the schedule of $\lambda$ across layers/rounds
$\nabla F_{k}(\theta)$ vector Full local gradient on client $k$
$g_{k}^{t}$ vector Stochastic gradient on client $k$ at round $t$
$g_{k,s}^{t}$ vector Stochastic gradient at local step $s\in\{0,\dots,\tau{-}1\}$ in round $t$
$L$ scalar Smoothness constant (Lipschitz gradient)
$\mu_{\text{sc}}$ scalar Strong convexity constant (when assumed in Sec. 4)
$\kappa$ scalar Condition number $L/\mu_{\text{sc}}$ (strongly convex case)
$\sigma^{2}$ scalar Upper bound on gradient variance
$\zeta_{k}$ scalar Gradient dissimilarity: $\sup_{\theta}\|\nabla F(\theta){-}\nabla F_{k}(\theta)\|^{2}$
$\zeta$ scalar Aggregate dissimilarity (e.g., $\sum_{k}\zeta_{k}$ or mean variant)
$\Delta_{k}$ scalar Local–global optimum gap $\|\theta_{k}^{\star}{-}\theta^{\star}\|^{2}$
$\mathcal{G}$ scalar Gradient moment bound used in multi-step analysis (Sec. 4)
$\mathrm{Beta}(\alpha_{\mathrm{B}},\alpha_{\mathrm{B}})$ dist. Beta distribution used to sample $\lambda$ in MixUp
Notes. (i) Aggregation-as-FedSGD yields $\boxed{\bar{\lambda}^{\,t}=1-\eta_{g}/\eta_{\ell}}$ and $\eta_{\mathrm{eff}}^{\,t}=(1{-}\bar{\lambda}^{\,t})\eta_{\ell}$. (ii) We reserve $\mu_{\text{sc}}$ for strong convexity; $\mu$ denotes the mixup scheduling factor if present in Methodology.
4 Theoretical Analysis
In this section we establish a clean theoretical view of pMixFed. We show that the update rule is equivalent to stochastic gradient descent (SGD) on the global objective with an effective step size governed by the mixup coefficients. This both clarifies stability conditions and explains how pMixFed mitigates catastrophic forgetting.
In FedSGD, the gradients are aggregated and the server will be update the global model according to the aggregated gradients. FedSGD is sometimes preferred over FedAvg due to its potentially faster convergence. However, it lacks robustness in heterogeneous environments. pMixFed leverages the faster convergence characteristics of FedSGD by incorporating early stopping mechanisms, facilitated by the use of mixup. As It’ll be demonstrated in Section 5, the mixup factor $\lambda$ functions analogously to an SGD update at the server, even though pMixFed aggregates model weights rather than gradients, similar to FedAvg. Section 4.2 provides a more detailed explanation of this mechanism.
4.1 Assumptions
Let $F(\theta):=\sum_{k=1}^{K}\Omega_{k}F_{k}(\theta)$ denote the global objective, where $\Omega_{k}=|D_{k}|/|D|$ and $\theta^{t}$ is the global model at round $t$. Each client $k$ runs $r$ local SGD steps with step size $\eta_{\ell}$, producing $\theta_{k}^{t+1}$. Mixup coefficients $\lambda_{k,i}^{t}\in[0,1]$ apply at layer $i$ of client $k$. For simplicity we define client-averaged $\lambda_{k}^{t}:=\sum_{i}w_{i}\lambda_{k,i}^{t}$ with $\sum_{i}w_{i}=1$, and population average $\bar{\lambda}^{t}:=\sum_{k}\Omega_{k}\lambda_{k}^{t}$. We make the following standard assumptions. These assumptions are standard in convergence analysis and ensure that the optimization process is well-behaved.
Assumption 1
(Unbiased Gradient Estimation with Bounded Variance.) The gradient estimate for local model updates is unbiased and has a bounded variance, i.e.,)
$\mathbb{E}[g_{k}^{t}]=\nabla F_{k}(\theta^{t})\quad\text{and}\quad\mathbb{E}\|g_{k}^{t}-\nabla F_{k}(\theta^{t})\|^{2}\leq\sigma^{2}.$ (4)
Assumption 2 (L-Smoothness)
Each $F_{k}$ has $L$-Lipschitz gradients, i.e.,
$\|\nabla F_{k}(x)-\nabla F_{k}(y)\|\leq L\|x-y\|$ (5)
Assumption 3 (Bounded Gradients)
The gradients at each client are bounded, i.e.,
$\mathbb{E}\|g_{k}^{t}\|^{2}\leq G^{2}$ (6)
4.2 Equivalence to FedSGD
In pMixFed, the server forms a convex combination between the previous global iterate and the locally updated models:
$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \sum_{k\in\mathcal{U}_{t}}\Omega_{k}\left(\lambda_{k}^{t}\,\theta^{t}+(1-\lambda_{k}^{t})\,\theta_{k}^{t+1}\right)\penalty 10000\ =\penalty 10000\ \bar{\lambda}^{\,t}\,\theta^{t}+(1-\bar{\lambda}^{\,t})\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\theta_{k}^{t+1}.$ (7)
With one local step from $\theta^{t}$, $\theta_{k}^{t+1}=\theta^{t}-\eta_{\ell}\,g_{k}^{t}$, (7) becomes
$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \theta^{t}-\underbrace{(1-\bar{\lambda}^{\,t})\,\eta_{\ell}}_{\displaystyle\eta_{\mathrm{eff}}^{\,t}}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}.$ (8)
Thus pMixFed is exactly centralized SGD on $F$ with effective step size
$\boxed{\qquad\eta_{\mathrm{eff}}^{\,t}\penalty 10000\ =\penalty 10000\ (1-\bar{\lambda}^{\,t})\,\eta_{\ell}.\qquad}$
Matching the FedSGD view.
If the server instead writes a FedSGD step with global step size $\eta_{g}>0$, i.e., $\theta^{t+1}=\theta^{t}-\eta_{g}\sum_{k}\Omega_{k}g_{k}^{t}$, then (8) shows the coefficients must satisfy:
Theorem 1 (Coefficient matching with FedSGD)
For any round $t$, the pMixFed update (7) similar with a FedSGD step of size $\eta_{g}$ if and only if
$\boxed{\qquad\bar{\lambda}^{\,t}\penalty 10000\ =\penalty 10000\ 1-\frac{\eta_{g}}{\eta_{\ell}}\,,\qquad 0\leq\frac{\eta_{g}}{\eta_{\ell}}\leq 1.\qquad}$
Proof Equate (8) with $\theta^{t+1}=\theta^{t}-\eta_{g}\sum_{k}\Omega_{k}g_{k}^{t}$ and match the coefficient of the aggregated gradient: $(1-\bar{\lambda}^{\,t})\eta_{\ell}=\eta_{g}$.
Interpretation.
Larger $\eta_{g}/\eta_{\ell}$ implies smaller $\bar{\lambda}^{\,t}$, i.e., less carry-over of $\theta^{t}$ and more trust in the new local updates. Standard smoothness suggests choosing $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$, equivalently $\bar{\lambda}^{\,t}\geq 1-\frac{1}{L\eta_{\ell}}$ whenever $L\eta_{\ell}>1$.
4.3 Convergence guarantees
The SGD view (8) directly yields standard optimization guarantees. We first state a one-step descent inequality, then derive the nonconvex and strongly-convex global results.
Lemma 2 (One-step descent)
Under (Asuumption1)–(Assumption2) and $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$,
$\mathbb{E}\big[F(\theta^{t+1})\big]\penalty 10000\ \leq\penalty 10000\ \mathbb{E}\big[F(\theta^{t})\big]\penalty 10000\ -\penalty 10000\ \frac{\eta_{\mathrm{eff}}^{\,t}}{2}\,\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\penalty 10000\ +\penalty 10000\ \frac{L\,(\eta_{\mathrm{eff}}^{\,t})^{2}}{2}\,\sigma^{2}.$
Proof By $L$-smoothness, $F(\theta^{t+1})\leq F(\theta^{t})+\langle\nabla F(\theta^{t}),\theta^{t+1}-\theta^{t}\rangle+\frac{L}{2}\|\theta^{t+1}-\theta^{t}\|^{2}$. Substitute (8), take conditional expectation using $\mathbb{E}[\sum_{k}\Omega_{k}g_{k}^{t}]=\nabla F(\theta^{t})$ and $\mathbb{E}\|\sum_{k}\Omega_{k}g_{k}^{t}\|^{2}\leq\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}$, and use $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$ to absorb the quadratic term in $\|\nabla F(\theta^{t})\|^{2}$.
Theorem 3 (Nonconvex rate)
Let $\eta_{\mathrm{eff}}^{\,t}\equiv\eta_{\mathrm{eff}}\leq 1/L$ be constant. Then for any $T\geq 1$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\penalty 10000\ \leq\penalty 10000\ \frac{2\,(F(\theta^{0})-F^{\star})}{T\,\eta_{\mathrm{eff}}}\penalty 10000\ +\penalty 10000\ L\,\eta_{\mathrm{eff}}\,\sigma^{2}.$
Proof Sum Lemma 2 over $t=0,\dots,T-1$ and telescope; divide by $T$. Please see appendix . 9
Theorem 4 (Strongly-convex case)
Under (A1)–(A3), with $\eta_{\mathrm{eff}}^{\,t}\equiv\eta_{\mathrm{eff}}\leq\min\{1/L,\,1/(2\mu)\}$,
$\mathbb{E}\big\|\theta^{t}-\theta^{\star}\big\|^{2}\penalty 10000\ \leq\penalty 10000\ (1-\mu\,\eta_{\mathrm{eff}})^{t}\,\big\|\theta^{0}-\theta^{\star}\big\|^{2}\penalty 10000\ +\penalty 10000\ \mathcal{O}\!\left(\frac{\sigma^{2}}{\mu}\,\eta_{\mathrm{eff}}\right).$
Proof Apply the standard strongly-convex SGD recursion with step $\eta_{\mathrm{eff}}$ to (8). Please see appendix . 9
4.4 Layer-wise mixing and forgetting control
Layer-wise coefficients $\lambda^{t}_{k,i}$ (with averaging weights $w_{i}$) produce the same effective step via $\bar{\lambda}^{\,t}=\sum_{k,i}\Omega_{k}\,w_{i}\,\lambda^{t}_{k,i}$ and $\eta_{\mathrm{eff}}^{\,t}=(1-\bar{\lambda}^{\,t})\eta_{\ell}$. The per-round drift satisfies
$\big\|\theta^{t+1}-\theta^{t}\big\|\penalty 10000\ =\penalty 10000\ \eta_{\mathrm{eff}}^{\,t}\,\Big\|\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}\Big\|.$
Hence larger $\bar{\lambda}^{\,t}$ explicitly caps the amount a single round can overwrite the global representation, limiting catastrophic forgetting, while smaller $\bar{\lambda}^{\,t}$ accelerates adaptation. Maintaining $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$ ensures stability irrespective of the layer schedule.
4.5 Multiple local steps per round
Suppose each selected client performs $\tau\geq 1$ local SGD steps from $\theta^{t}$: $\theta_{k}^{t+1}\penalty 10000\ =\penalty 10000\ \theta^{t}-\eta_{\ell}\sum_{s=0}^{\tau-1}g_{k,s}^{t}.$ Then
$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \theta^{t}\penalty 10000\ -\penalty 10000\ (1-\bar{\lambda}^{\,t})\,\eta_{\ell}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\sum_{s=0}^{\tau-1}g_{k,s}^{t}\penalty 10000\ \equiv\penalty 10000\ \theta^{t}\penalty 10000\ -\penalty 10000\ \eta_{\mathrm{eff},\tau}^{\,t}\,\widehat{g}^{\,t},\qquad\eta_{\mathrm{eff},\tau}^{\,t}:=(1-\bar{\lambda}^{\,t})\,\eta_{\ell}\,\tau,$
where $\widehat{g}^{\,t}$ is the average of the $\tau$ local gradients. Under (Assumption1)–(Assumption2) and bounded second moments, standard local-SGD arguments give
$\big\|\mathbb{E}[\widehat{g}^{\,t}]-\nabla F(\theta^{t})\big\|\penalty 10000\ \leq\penalty 10000\ c_{1}\,L\,\eta_{\ell}\,(\tau-1)\,G,$
for a constant $c_{1}$ and gradient-moment bound $G$. Consequently, pMixFed behaves like SGD with step $\eta_{\mathrm{eff},\tau}^{\,t}$ up to a bias of order $(1-\bar{\lambda}^{\,t})\,L\,\eta_{\ell}^{2}\,\tau(\tau-1)$. Choosing $\bar{\lambda}^{\,t}$ so that $\eta_{\mathrm{eff},\tau}^{\,t}\leq 1/L$ recovers the guarantees of Theorems 34 with an additional (controlled) residual proportional to $L\,\eta_{\ell}^{2}\,\tau(\tau-1)$.
Remark2. In practice, this relationship suggests that by tuning $\lambda_{k}$, we can control the influence of the global model versus the local models in the aggregation process, similar to adjusting the learning rate in FedSGD. This is particularly beneficial in heterogeneous environments where clients may have varying data distributions.
Our theoretical analysis indicates that the mixup coefficient $\lambda_{k}$ in pMixFed plays a role analogous to the learning rate in FedSGD. This equivalence provides a deeper understanding of how pMixFed leverages the strengths of FedSGD while mitigating its weaknesses in heterogeneous settings. By appropriately choosing $\lambda_{k}$, pMixFed can achieve faster convergence and improved robustness. The pMixFed algorithm combines the advantages of FedSGD and FedAvg by aggregating model weights rather than gradients, while still ensuring faster convergence even in heterogeneous data settings due to it’s adaptation ability. The incorporation of the mixup mechanism enhances stability, providing faster convergence rates compared to FedSGD, particularly in non-convex settings
5 Methodology
In this section, we present our FL approach, which incorporates layer-wise Mixup in the feature space. We begin by briefly defining the Mixup technique, followed by a detailed description of the proposed algorithm and its integration into the federated setting. We also discuss the rationale behind this design and explain how it facilitates improved personalization across clients.
@@ 5.1 Mixup
Mixup is a data augmentation technique for enhancing model generalization zhang2017mixup based on learning to generalize on linear combinations of training examples. Variations of Mixup have consistently excelled in vision tasks, contributing to improved robustness, generalization, and adversarial privacy. Mixup creates augmented samples as:
$\begin{split}\bar{x}=\lambda.x_{i}+(1-\lambda).x_{j}\\ \bar{y}=\lambda.y_{i}+(1-\lambda).y_{j},\end{split}$ (9)
where $x_{i}$ and $x_{j}$ are two input samples, $y_{i}$ and $y_{j}$ are the corresponding labels, $\lambda$, $\lambda\sim\text{Beta}(\alpha,\alpha),\lambda\in[0,1]$, is the degree of interpolation between the two samples. Mixup relates data points belonging to different classes which has been shown to be successful in mitigating overfitting and improving model generalization verma2019manifold; guo2019mixup; zhang2020does. There are many variants of Mixup that has been developed to address specific challenges and enhance its effectiveness. For instance, AlignMixup venkataramanan2022alignmixup improves local spatial alignment by introducing transformations that better preserve semantic consistency between input pairs. Manifold Mixup verma2019manifold extends the concept to hidden layer representations, acting as a powerful regularization technique by training deep neural networks (DNNs) on linear combinations of intermediate features. CutMix yun2019cutmix replaces patches between images, blending visual information while retaining spatial structure. Remix chou2020remix addresses class imbalance by assigning higher weights to minority classes during the mixing process, enhancing the robustness of the trained model on imbalanced datasets. AdaMix guo2019mixup dynamically optimizes the mixing distributions, reducing overlaps and improving training efficiency. This linear interpolation also serves as a regularization technique that shapes smoother decision boundaries, thereby enhancing the ability of a trained model to generalize to unseen data. Mixup also increases robustness against adversarial attacks zhang2020does; beckham2019adversarial; and improves performance against noise, corrupted labels, and uncertainty as it relaxes the dependency on specific information guo2019mixup.
Figure 2: Workflow of pMixFed: Mixup is used in two stages. 1-Broadcasting: when transferring knowledge to local models, the frozen personalized model $L_{k}^{(t)}$ is mixed up with global model $G^{(t)}$ according to the adaptive mix factor $\mu_{k}^{(t)}$ which determines layer-wise mixup degree $\lambda_{i}$ for layer $i$. 2-Aggregation: The updated global model $G^{(t+1)}$ is generated through applying Mixup between the updated local model $L^{(t+1)}$ and the current global model $G^{\prime(t)}$ state.
@@
Our goal is to leverage the well-established benefits of Mixup in the context of personalized FL. While Mixup has previously been employed in FL frameworks, such as XORMixup shin2020xor, FEDMIX yoon2021fedmix, and FedMix wicaksana2022fedmix, prior studies have primarily focused on using Mixup for data augmentation or data averaging. We propose pMixFed by integrating Mixup on the model parameter space, rather than using it on the feature space. We apply Mixup between the parameters of the global and the local models in a layer-wise manner for more customized and adaptive PFL. Our approach eliminates the need for static and rigid partitioning strategies. Specifically, during both the broadcasting and aggregation stages of our partial PFL framework, we generate mixed model weights using an interpolation strategy which is illustrated in Figure 2. Mixup offers the flexibility in combining models by introducing a mix degree for each layer $\lambda_{i}$, which changes gradually according to $\mu$, i.e., the mix factor. Parameter $\mu$ is also updated adaptively in each communication round and for each client according to the test accuracy of the global and the local model during the evaluation phase of FL. $\mu$ is computed as follows:
$\mu_{k}^{t}=1-\frac{1}{1+e^{-\delta(Acc^{b}-1)}}$ (10)
$\delta=(t/T)$, where $t$, and $T$ are the current communication round, and the overall number of communication rounds respectively. Also, $Acc$ calculated as: $Acc=(Acc_{k}^{t}/Acc_{overall}^{(t-1)})$ in the broadcasting phase and average test accuracy of the previous global model $\left(Acc=G^{t}(x,\theta^{t}),y\right)$ on all local test sets $x=\{x_{1},x_{2},\ldots,x_{K}\}$. , in the aggregation stage. $b$ is the offset parameter for sigmoid function which is set to $2$ after several experiments. More detailed discussion on parameter $\mu$’s rule of update is discussed in section 5 and the reason behind this formulation is discussed in the Appendix Sec. 2.
@@
$\begin{split}\text{Broadcasting Stage:}\quad\quad\lambda_{i}=\begin{cases}1&\lambda_{i}>1\\ \mu*(n-i)&\lambda_{i}\leq 1\end{cases}\\ \text{Aggregation Stage:}\quad\quad\lambda_{i}=\begin{cases}0&\lambda_{i}\leq 0\\ 1-(i*\mu)&\lambda_{i}>0,\end{cases}\end{split}$ (11)